OpenStudy (anonymous):

Hi I'm doing a quiz but I got stuck on a problem and I got to 14 x +3 = 3y -10 and I kinda forgot how to solve from there can some one help me so I can find the answer.

OpenStudy (snowsurf):

Looking at this equation 14 x +3 = 3y -10 what terms can you combine?

OpenStudy (snowsurf):

Which look alike so that you can combine them by adding or subtracting.

OpenStudy (anonymous):

sorry 14x not multiply

OpenStudy (triciaal):

you cannot solve a linear equation with 2 unknown variables

OpenStudy (anonymous):

can you help me from the beginning of the problem

OpenStudy (triciaal):

can you post the original question?

OpenStudy (anonymous):

(8x -8 = x-5) = (3y - 6 = 7x +4)

OpenStudy (anonymous):

solve for x and y

OpenStudy (triciaal):

when you solve you are usually finding the value for an unknown

OpenStudy (anonymous):

and I'm solving for x and y

OpenStudy (anonymous):

which would be my unkown

OpenStudy (triciaal):